Solve using the multiplication principle. Don't forget to perform a check. -3/4x=-5/6

Answer by checkley77(12844)

3/4x=-5/6
3X/4=-5/6
3X*6=4*-5
18X=-20
X=-20/18
X=-10/9 ANS.
PROOF:
3/4*-10/9=-5/6
-30/36=-5/6
-5/6=-5/6